OpenStudy (chosenmatt):

What is the value of the expression –5x4 when x = –2? A. –80 B. 80 C. –40 D. 40

OpenStudy (chosenmatt):

@iGreen

OpenStudy (chosenmatt):

@sammixboo

OpenStudy (chosenmatt):

@mathstudent55

OpenStudy (chosenmatt):

@marylou004

OpenStudy (danjs):

i think you typed that wrong. looks just like -5 times 4

OpenStudy (igreen):

x = -2 \(-5x^{4}\) \(-5(-2)^4\) Okay, the exponent is on -2. So multiply \(-2 \times -2\) first.

sammixboo (sammixboo):

Does it look like this? \(-5x^4\)?

OpenStudy (chosenmatt):

high 4 lol

OpenStudy (igreen):

@sammixboo I'm pretty sure, his last question was about exponents.

OpenStudy (chosenmatt):

yes!!!

OpenStudy (igreen):

Yep, that gives us -5(4). So can you multiply that?

sammixboo (sammixboo):

OK so if x = -2, plug in the value(s) :) \(-5x^4\) \(-5\times-2^4\) So first do the exponent, \(-2^4\) = -2 x -2 x -2 x -2

sammixboo (sammixboo):

So what is -2 x -2 x -2 x -2?

OpenStudy (uri):



OpenStudy (uri):

Solve it @chosenmatt

OpenStudy (igreen):

Ugh, stupid OS, my comments are lagging -.-

sammixboo (sammixboo):

Um, I thought it was \(-2^4\), not \(-2^2\)

OpenStudy (igreen):

Oops, that's -2 x -2 x -2 x -2 >.> My brain feels like sludge today..

OpenStudy (igreen):

I'm signing off, see ya guys!

OpenStudy (chosenmatt):

-20?

OpenStudy (uri):

Well you should know about Pemdas,Solving exponents comes first SO, -2^4 and then multiply the answer by -5

OpenStudy (igreen):

@chosenmatt It's going to be positive.

OpenStudy (chosenmatt):

40

sammixboo (sammixboo):

Mm try again :)

OpenStudy (uri):

-2^4 =-16

OpenStudy (chosenmatt):

yeah ositive 40

OpenStudy (chosenmatt):

POSITIVE LOL

OpenStudy (chosenmatt):

answer choices is A. –80 B. 80 C. –40 D. 40

OpenStudy (chosenmatt):

@sammixboo

sammixboo (sammixboo):

Can you tell me what is -2 x -2 x -2 x -2 is first?

OpenStudy (chosenmatt):

is it negative 40? @sammixboo

sammixboo (sammixboo):

Nope, the answer is positive

OpenStudy (uri):

-16*-5 =